OpenStudy (anonymous):

simplify a^2-ab/a^2+ab

OpenStudy (anonymous):

(a - b)² = a² -2ab + b² (a + b)² = a² + 2ab + b² So to get from (a - b)² to (a + b)² you have to add 4ab so: (a - b)² + 4ab = (a + b)² Or: (a - b)² = (a + b)² - 4ab

OpenStudy (anonymous):

i was so gonna say that ^^